OpenStudy (anonymous):

OK so pre-calc question: solve the triangle A=50, b=13, c= 6 Am I doing this correctly and if I am can you guide me to the next process, however if I've goofed up help me to the right path cos A = b^2 + c^2 - a^2 / 2bc cos 50 = 13^2 + 6^2 - a^2 / 2(13)(6) cos 50 = 205 - a^2 / 156 ?????

OpenStudy (anonymous):

You are almost there. Next step is to set the equation a^2/156=205-cos 50 then you get a^2=205-cos 50 times 156 From this point, I use calculator to find a. Good Luck